OpenStudy (anonymous):

bailey's banquet hall has 16 tables that seat 8 people each and 1 table that seats 12 people.What is the total number of seats in the banquet hall. a.96 b.128 c.192 d.140

OpenStudy (anonymous):

Using simple multiplication: Since 16 tables seat 8 people then 16*8=128. With the addition of 1 table that seats 12 people then 1*12=12. Add them together and you get 128+12=140